# A Versatility Measure for Parametric Risk Models

### v1

@@ 7 Conclusion
  • The versatility measure of a 2-parameter distribution often falls between the measures of two 1-parameter special cases (with each of the parameters fixed in turn), suggesting that the characteristics contributed by the two parameters individually are “averaged” when taken together.
@@
  • The versatility measures of certain distributions (i.e., $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$) are considerably lower than other measures, raising serious concerns about idiosyncratic behavior.
Overall, we would argue that the proposed versatility measure may be employed to assess the potential usefulness of probability distributions a priori by accounting for and mitigating problems of idiosyncratic overfitting. Although our analysis does not propose minimum cutoff values for model-selection purposes, the versatility measure can be used as an effective guide for identifying a range of acceptable values based on the measures arising from distribution families found to be useful in practice. For example, the relatively low measures associated with the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ family should not necessarily disqualify this commonly used model, but rather suggest a lower region for reasonable versatility levels. On the other hand, the substantially lower measures computed for the $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$ distributions likely suggest these models should be avoided.
Much further work remains to be done in this area, particularly involving the analysis of marginal contributions to versatility of additional parameters for $k\geq 2$. In this regard, it seems likely that the most useful potential applications of the proposed versatility measure lie in the investigation and comparison of probability distributions within multiple-parameter families; for example, the Generalized Beta, Burr, Pareto, and Pearson families increasingly used to model severities in risk analysis.

### v3

@@ 7 Conclusion
  • The versatility measure of a 2-parameter distribution often, but not always, falls between the measures of two 1-parameter special cases (with each of the parameters fixed in turn), suggesting that the characteristics contributed by the two parameters individually tend to be “averaged” when taken together.
@@
  • The versatility measures of certain distributions (i.e., Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$, $\textrm{Lognormal}\left(\ln\left(\nu\right),\sigma=1\right)$, Negative $\textrm{Negative Binomial}\left(r,\tfrac{m}{m+1}=\tfrac{1}{2}\right)$, and $\textrm{Waring}\left(\alpha=1,\vartheta\right)$) are noticeably lower than those of other distributions, raising serious concerns about idiosyncratic behavior.
Finally, by comparing the Pareto $\textrm{Pareto 2}\left(\alpha,\vartheta\right)$ model with the more versatile Inverse $\textrm{Inverse Gamma}\left(r,\lambda\right)$ alternative, we showed how the proposed versatility measure may be employed to assess the usefulness of probability distributions a priori, effectively recognizing and mitigating problems of idiosyncratic overfitting. Although our analysis does not propose minimum cutoff values for model-selection purposes, we believe that, by observing large numbers of versatility measures in applied work, it will be possible to establish reasonable empirical guidelines for acceptable ranges of $\mathcal{V}\left(F_{0}\right)$.
Much further research remains to be done on this topic, particularly involving the analysis of marginal contributions to versatility of additional parameters for $k\geq 2$. In this regard, it seems likely that the most useful potential applications of the proposed versatility measure lie in the investigation and comparison of probability distributions within: (i) multiple-parameter continuous families, such as the Generalized Beta, Burr, Pareto, and Pearson distributions used to model severities; and (ii) discrete families modified to fit individual parameters to specific values within the sample space (especially $x=0$ or $1$), such as the Zero-Inflated Poisson and Negative Binomial models for frequencies.